OpenStudy (anonymous):

Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3)

OpenStudy (anonymous):

could someone explain this?

OpenStudy (anonymous):

Plug in the numbers for x and y and see if you get the same answer it eguals.

OpenStudy (anonymous):

That's what I thought thanks :)

OpenStudy (anonymous):

\[2*4-5*1=3\] so your answer is A.

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

Wait no Sorry.

OpenStudy (anonymous):

Yeah it is A. I forgot to do the 2nd part. But it still is the same match. So, I would definetly choose A.

OpenStudy (anonymous):

I'l tell you if it was right after the test. Could you help me with two more? @cherry18

OpenStudy (anonymous):

Sure. If i can. Btw your not suppose to do test questions. But i don't care so post away.

OpenStudy (imtiaz7):

A is your right answer

OpenStudy (anonymous):

You don't have to give an answer, If you could just explain how to solve it, that would be great :)

OpenStudy (anonymous):

Ok. I am fine either way. Have anymore questions to answer/explain?

OpenStudy (imtiaz7):

so here i am going to tell you that you first need to multiply second equation by 2 we got: 2x-6y=2 and then subtract it by equation 1 we got: 2x-5y=3 so we get x=4 now put x's value in equation 1 we got: 2(4)-5y=3 8-5y=3 -5y=-5 y=1

OpenStudy (imtiaz7):

do you get it?